Dragbin, Inc., is trying to determine its cost of debt. The firm has a debt issue outstanding with 16 years to maturity that is quoted at 106 percent of face value. The issue makes semiannual payments and has an embedded cost of 6 percent annually.

What is the company’s pretax cost of debt?(Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

If the tax rate is 35 percent, what is the aftertax cost of debt? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1
We need to use Funtion RATE to find the Yield of Bond
NPER 16*2 =Number of Payments 32
PMT =6%/2*100 =Semi annual Coupen Payment 3
PV=Value of Bond -106
FV =future Value of Bond 100
=RATE(32,3,-106,100)*2 5.43%
Cost of Debt =5.43%
After tax Cost of debt =5.43%*(1-.35) 3.53%